1) If the government wants to reduce unemployment, government purchases should be ________ and/or taxes should be ________.

a) increased; decreased

b) increased; increased

c) decreased; decreased

d) decreased; increased

2) The President of Vulcan hires you as an economic consultant. He is concerned that the output level in Vulcan is too high and that this will cause prices to rise. He feels that it is necessary to reduce output by $10 billion. He tells you that the MPC in Vulcan is 0.6. Which of the following would be the best advice to give to the Vulcan president?

a) reduce government purchases by $10 billion

b) increase taxes by $2.5 billion

c) increase taxes by $10 billion

d) reduce government purchases by $4 billion
Answer #1

1. Ans: a) increased; decreased

Explanation:

Money supply in the economy increases through government spending by multiplier effect. Investment opportunity will increase. As a result unemployment will be reduced. Reduction in tax rate also reduces unemployment

2. Ans: d) reduce government purchases by $4 billion

Explanation:

Multiplier = 1 / (1 - MPC) = 1 / ( 1 - 0.6) = 1/ 0.4 = 2.5

Reduction in government purchase = $10 billion / 2.5 = $4 billion